OpenStudy (anonymous):

The volume of gas varies directly with temperature and inversely with pressure. Volume is 100 m^3 when temperature is 1508 degrees and pressure is 15 lb/cm^2. What is the volume when the temperature is 2508 degrees and the pressure is 20 lb/cm^2....

OpenStudy (anonymous):

anyone explain?

OpenStudy (anonymous):

Ideal Gas Law PV=nRT

OpenStudy (anonymous):

P is pressure, V is volume, n, is the number of molecules, R is a constant, and T is temperature. n and R will not change, so set up a ratio with the other values.

OpenStudy (anonymous):

hmm?

OpenStudy (zpupster):

i thought i answered this yesterday: Gas Equation: PiVi/Ti = PfVf/Tf PV= nrT is the ideal Gas LAw but for this better the given formula above stands for i= initial f= final what we know Pi = 15 lb/cm^2. Vi = 100 m3 Ti = 1508K Pf = 20 lb/cm^2. Tf = 2508 K what we want is the final Volume Vf= ???? lets plug these into the given equation: PiVi/Ti = PfVf/Tf (15)(100) / 1508 = (20)(Vf) / 2508 now for the math: we divide both sides by 20 and multiply by 2508 20 and 2508 cancel on left side and we have (15)(100)(2508) /(1508)(20) = Vf put this in your calculator i did not put the units --but they cancel leaving =Vf in m3

OpenStudy (anonymous):

49,893.8?

OpenStudy (zpupster):

124.734